40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 1892     Entry time: Wed Jan 13 17:52:35 2021
Author: Paco 
Type: DailyProgress 
Category: Optics 
Subject: OPO cavity resonance 

Observed first resonant transmitted (& reflected) light from the DOPO cavity; the PZT scan was centered at 31 V, at 2 Hz, with an amp. of 1.5 Vpp. To get there, revisited the path's alignment upstream to the last mirror (before the last lens), removing, inspecting, and reinstalling each component. After this, I used the camera at the end of the optical path as a "pinhole" (beam center placeholder) and after inserting each element (mirrors / crystal) checked carefully that the beam was landing straight. Then, patiently scanned various knobs (mirror mounts X/Y/XY, crystal manually) until HOM started resonating. After a bit of further alignment managed to see transmission dips in the FI pickoff. Below are two photos illustrating the current state (way more optimization is needed), as well as the setup viewed from one side (for the scope picture, purple is the ramp, yellow is cavity reflection, green is cavity transmission). Will keep optimizing in the couple next days, all at low power first, and then start cranking the power up to factor in any thermal effects into the optimized cavity.

Attachment 1: dopo_first_resonance.jpg  3.674 MB  | Hide | Hide all | Show all
dopo_first_resonance.jpg
Attachment 2: dopo_sideview.jpg  3.577 MB  | Show | Hide all | Show all
ELOG V3.1.3-